OpenStudy (anonymous):

y=20+5(25-(1/2)x) +(1/2).(25-(1/2)x)^2 it turns in to y=(1/8)x^2 -15x + 457,5 i have no idea how that happened, someone mind explaining? thnx!

OpenStudy (anonymous):

\[y=20+5(25-\frac{ 1 }{ 2 }x)+(\frac{ 1 }{ 2 })(25-\frac{ 1 }{ 2}x)^{2}\] and now use foil, remember, dont put it to decimals and \[(25-\frac{ 1 }{ 2 }x)^2=(25-\frac{ 1 }{ 2 }x)(25-\frac{ 1 }{ 2 }x)\]

OpenStudy (anonymous):

try to simplify it now.

OpenStudy (anonymous):

i'm sorry i'm lost, where did the (1/2) go and how are there 2 (25-(1/2)x)?

OpenStudy (anonymous):

no i didnt give you the answer i was just saying that (25-(1/2)x)^2 is being multiplied by itself. so first expand that out and then simplify

OpenStudy (anonymous):

oh i see, now it make sense, thank you very much!

OpenStudy (anonymous):

can you do that for every bracket powered to the 2nd?

OpenStudy (anonymous):

yes, you can do that with every power greater than 1.